OpenStudy (anonymous):

What is the sum of the arithmetic series? Can someone explain this it's confusing me quite a bit. (Will post equation)

OpenStudy (anonymous):

\[\sum_{t = 1}^{18} (3t - 4)\]

OpenStudy (anonymous):

The formula for the first n terms of an arithmetic sequence, starting with n = 1, is: \[\sum_{i=1}^{n}a_i = (n/2) (a_1+a_n)\] So plug-in and solve. \[(18/2)(-1+50)= 9*49=441\]

OpenStudy (anonymous):

Thanks, I get that now.. Do you know about the sum of finite geometric series?

OpenStudy (anonymous):

Yea, one sec.

OpenStudy (anonymous):

\[\sum_{k=0}^{n-1}ar^k = a * ((1-r^n)/(1-r))\]

OpenStudy (anonymous):

For\[\sum_{ = 0}^{15} 2(1/2)^{x}\] Does that mean: \[2(\frac{ 1 - 1/2^{15} }{ 1 - 12 }) \] Is that the right way to set it up?

OpenStudy (anonymous):

I would think so. That makes sense.

OpenStudy (anonymous):

Thanks a lot!